OpenStudy (mtwizzle420):

Simplify the rational expression. @misty1212

OpenStudy (misty1212):

HI!1 (again) is this the same type problem?

OpenStudy (mtwizzle420):



OpenStudy (misty1212):

\[125=25\times 3\] so \[\sqrt{125}=5\sqrt3\]

OpenStudy (misty1212):

and \(\sqrt{h^4}=h^2\) final answer \[\huge 5h^2\sqrt3\]